I'm having a hard time solving the following question: Naima’s backyard is the shape of a rectangle. The length is 5 m less than twice the width, and the perimeter is 80 m. Find the dimensions of the yard
Found 2 solutions by stanbon, solver91311:
Answer by stanbon(75887)

Naima’s backyard is the shape of a rectangle. The length is 5 m less than twice the width, and the perimeter is 80 m. Find the dimensions of the yard
-----
Draw the picture.
P = 2L + 2W = 80m
L + W = 40m
---
L = 2W-5
---
Substitute and solve for "W".
2W-5+W = 40
3W = 45
Width = 15m
----
Solve for "L":
L = 2W-5
L = 2*15-5
Length = 25m
======================
Cheers,
Stan H.
=================

Answer by solver91311(24713)



If a rectangle with perimeter has a length that is units longer (or shorter if ) than times the width, ,

Then knowing that the Perimeter is given by:

,

make the substitution:



Then solve for







For your problem: , , and . Do the arithmetic.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism